# Large Language Model

### v1

@@ 5.1 Experiments Setup
Datasets. N-ImageNet [23] is the event camera version of ImageNet and the largest event camera dataset, which contains almost 1.2 million event streams and 1,000 classes. HARDVS [40] is a recently released dataset for event-based action recognition, currently having the largest action classes, namely 107,646 recordings for 300 action categories. DSEC [16] (Driving Stereo Event Camera) dataset is an autonomous driving dataset with a particular focus on event-based vision. We also utilize N-Caltech101 [31] for evaluating model zero-shot performance and further few-shot testing. Similar to N-ImageNet construction, N-Caltech101 contains 8,246 samples from 101 classes. More dataset details can be found in the appendix.
@@
Dataset Preprocessing. Overall processed dataset information can be referred to Table 1. After we annotate these dataset by Section 3, we split the image/video-event-text pairs from N-ImageNet [23], HARDVS [40], N-Caltech101 [31] into the corresponding training sets and validation sets according to their categories. We use the whole DSEC training set by segmenting its scene video into multiple frame groups. The categories in each split do not overlap with each other in the same dataset for zero-shot evaluation. Few-shot datasets setting can be found in the appendix.
@@
Figure 4: Qualitative results about zero-shot caption generation by utilizing DSEC [16] origin val set. Red color denotes model wrong description and orange color denotes the accurate object description.
5.2 Caption Generation
Zero-shot Evaluation. To evaluate the event-based understanding performance of our proposed framework, we conduct comparisons with the selected SOTA MLLM baselines, including image-based [38, 30, 29] and video-based MLLMs [28, 9]. As shown in Table 2, we can conclude that the video-based MLLMs can generate more high-quality descriptions for event data, which is determined by the stream properties of event data. Finally, experiment results demonstrate the effectiveness of EventVL to describe the event data. Our proposed EventVL significantly achieves 2.603 and 0.573 CIDEr in HARDVS and N-ImageNet respectively, outperforming existing works by a large margin, showing our advantages in event understanding ability. Furthermore, our model parameters are only 2.3B, proving its effectiveness, which is easy to deploy on the edge device.
Few-shot Evaluation. We also explore the impact of the training data scale on our event-based understanding performance. As shown in Table 3, the performance is closely related to the size of the dataset, which basically conforms to the principle of Scaling Law [10]. When $n=5$, the model learns more class invariant knowledge, which is beneficial for making more comprehensive perceptions. We also conclude that in a little data ($n\leq 10$), our model can get a significant promotion in event analysis, which verifies our design superiority. However, the improvement of event-based understanding ability is gradually slowing down when $n\geq 20$, because event-based knowledge learning has approached its limits in the limited parameters.
5.3 Qualitative results
@@
We perform qualitative experiments on various scenario val sets with the zero-shot setting and utilize VideoLLaMA2 [9] to compare with our proposed EventVL. With only event data as input, as shown in Figure 3, our model can generate accurate descriptions for these event-based represenation, while VideoLLaMA2 always generate some useless text, like the denoted underline text “around in a field of dots”. The reason is that current MLLMs training data do not contain the event-based representation, which causes falling in explicit recognition wrong for factual hallucinations. In addition, our model can easily perceive objects and actions in hard scenes such as low light and blurry scenes, and provide comprehensive descriptions. Like the example in the upper left corner of HARDVS cases, our EventVL can not only describe this motion but also infer the colour of clothing worn by humans. Next, we also investigate the long description generation ability of EventVL by evaluating the drive scene. As shown in Figure 4, our proposed EventVL can generate a more precise description for the traffic scene and reasonable prediction ability like the second case “Traffic lights are green” in orange colour. However, VideoLLaMA2 easily has hallucinations when seeing these representations like the accident description in red color. Qualitative results show our advantage in understanding event-based presentation. Results about multiple rounds of dialogue and failure cases are also shown in the appendix.

### v2

@@ V-B Datasets
Datasets. N-ImageNet [15] is the event camera version of ImageNet and the largest event camera dataset, which contains almost 1.2 million event streams and 1,000 classes. HARDVS [16] is a recently released dataset for event-based action recognition, currently having the largest action classes, namely 107,646 recordings for 300 action categories. DSEC [25] (Driving Stereo Event Camera) dataset is an autonomous driving dataset with a particular focus on event-based vision. We also utilize N-Caltech101 [26] for evaluating model zero-shot performance and further few-shot testing. Similar to N-ImageNet construction, N-Caltech101 contains 8,246 samples from 101 classes.
@@
Processing Details Overall processed dataset information can be referred to Table I. In each annotation round, we randomly sample a problem from the corresponding list. For N-ImageNet, we use the entire event-pairs dataset for training, without utilizing its variants. For N-Caltech101, the dataset is only used to validate the effectiveness of our proposed EventVL. For HARDVS, we employ its red-blue color map representation rather than the raw event data to improve efficiency. In each category, we uniformly sample 14 frames from each video for post-processing. If a video contains fewer than 14 frames, we use the entire video. For DSEC, we sample 14 frames from each segment to ensure training efficiency. After we annotate these dataset, we split the image/video-event-text pairs from N-ImageNet, HARDVS, N-Caltech101 into the corresponding training sets and validation sets according to their categories. We use the whole DSEC training set by segmenting its scene video into multiple frame groups. The categories in each split do not overlap with each other in the same dataset for zero-shot evaluation.
V-C Implementation
Similar to event-based VLM [7, 12, 22], we process each event stream as the event-based frame, and input them into diverse MLLMs, including the normal image-based MLLM [19, 18, 13] and Video-based MLLM [27, 28]. Then, to further evaluate event-based MLLM comprehensively, we select InternVL [17] as the base model with additional event encoder copied from image encoder, Then we utilize the proposed datasets to finetune its text decoder, which named as InternVL-E.
Evaluate Metrices. Similar to previous works [29, 30], we select the Bleu, METEOR, ROUGE-L, CIDEr and VQA [31] for evaluation, which measure precision and contextual understanding in responding to specific queries. For VQA, we design well-structured prompts. We then leverage GPT for quality assessment using a scoring scale from 1 to 5 and establish a comprehensive evaluation benchmark based on the EventVL-QA datasets. This framework enables a rigorous and systematic evaluation of the proposed EventVL.
Figure 2: Qualitative results about zero-shot caption generation by utilizing N-ImageNet [15] and HARDVS [16] val set from Event-Base dataset. Red color denotes model wrong description and orange color denotes the accurate object description. Underline text means redundant description. It shows that EventVL can easily distinguish and analyze the vague objects imaging in the complex scenes like low light or motion-blur based on event streams.
@@ V-C Implementation
Figure 3: Qualitative results about zero-shot caption generation by utilizing DSEC [25] origin val set from Event-Base dataset. Red color denotes model wrong description and orange color denotes the accurate object description. It shows that EventVL can easily understand the complicated traffic scenes like low light and multiple objects based on event streams.
V-D Comparison with SOTAs
In following experiments, we only consider the event data as input for accurately evaluating the model’s event-based understanding ability.
Zero-shot Captioning. As shown in Table III, we can conclude that video-based MLLMs generate higher-quality descriptions for event data, primarily due to the temporal properties of event data. Additionally, the performance of InternVL-E highlights the significant modal information gap between event flows and conventional images, motivating further research on adapting existing MLLMs to event data. In Table III, our proposed EventVL achieves 2.619 and 0.599 CIDEr for HARDVS and N-Caltech, respectively, significantly outperforming existing methods. This demonstrates our model’s superior event understanding ability. With only2.3B parameters, our model proves highly effective and easily deployable on edge devices.
Question Answering. We also conduct the detailed QA experiments to evaluate the event-based analysis and understanding ability as shown in Table IV. For N-ImageNet, EventVL achieves a VQA score of 4.13, outperforming InternVL-E (3.34) and significantly surpassing Video-LLaVA (2.87). This indicates that EventVL effectively captures and understands event information in N-ImageNet, despite having far fewer parameters than Video-LLaVA. Moreover, the superior performance on HARDVS demonstrates the robustness of EventVL in handling more complex and dynamic event question-answering scenarios compared to other methods.
TABLE IV: Quantitative performance of VQA on different datasets.
Dataset Model #param VQA
N-ImageNet [15] Video-LLaVA [27] 7B 2.87
InternVL-E [13] 2.3B 3.34
EventVL (Ours) 2.3B 4.13
HARDVS [16] Video-LLaVA [27] 7B 3.23
InternVL-E [17] 2.3B 3.73
EventVL (Ours) 2.3B 4.26
V-E Qualitative results
@@
We perform qualitative experiments on various scenario val sets with the zero-shot setting and utilize VideoLLaMA2 [28] to compare with our proposed EventVL. With only event data as input, as shown in Figure 2, our model can generate accurate descriptions for these event-based representation, while VideoLLaMA2 always generate some useless text, like the denoted underline text “around in a field of dots”. The reason is that current MLLMs training data do not contain the event-based representation, which causes falling in explicit recognition wrong for factual hallucinations. In addition, our model can easily perceive objects and actions in hard scenes such as low light and blurry scenes, and provide comprehensive descriptions. Like the example in the upper left corner of HARDVS cases, our EventVL can not only describe this motion but also infer the colour of clothing worn by humans. Next, we also investigate the long description generation ability of EventVL by evaluating the drive scene. As shown in Figure 3, our proposed EventVL can generate a more precise description for the traffic scene and reasonable prediction ability like the second case “Traffic lights are green” in orange colour. However, VideoLLaMA2 easily has hallucinations when seeing these representations like the accident description in red color. Qualitative results show our advantage in comprehensive understanding event-based representation .
TABLE V: Ablation study of the proposed Event Spatiotemporal Representation (ESR). “HTS” denotes the Hierarchical Temporal Split and “ASS” denotes the Adaptive Spatial Split.
HTS ASS N-ImageNet HARDVS
ROUGE-L CIDEr ROUGE-L CIDEr
- - 0.352 0.806 0.493 2.274
- 0.384 0.833 0.615 2.512
- 0.377 0.821 0.611 2.478
0.397 0.853 0.639 2.619
V-F Multi-round Dialogue
As illustrated in Figure 4, we present several examples showcasing the application of our proposed EventVL framework in multi-round dialogue using event data. These examples demonstrate that EventVL not only understands the comprehensive semantics within event data but also captures finer object details, like describing “a clear blue sky”.
Figure 4: Results about multi-round dialogue by utilizing N-ImageNet [15] val set from Event-QA dataset. Blue color denotes the detailed description.
V-G Ablation Study
In this section, we utilize the zero-shot caption generation evaluation setting to perform the comprehensive ablation studies for investigating each effective module.
Event Representation. As shown in Table V, we investigate different impact of each module proposed in Event Spatiotemporal Representation. Enabling ASS alone improves performance across all metrics, which suggests that ASS enhances spatial feature representation which improves visual understanding and textual alignment. Besides, enabling HTS alone also yields improvements, indicating that HTS models temporal dependencies and captures event stream dynamics more comprehensively. Finally, the results show the ESR effectiveness in improving event-based multimodal learning and reinforce the importance of spatiotemporal representations for complex event-language tasks.
TABLE VI: Different $n_{\max}$ settings for Dynamic Semantic Alignment. Note that we default the $n_{\min}=1$.
$n_{\max}$ N-ImageNet HARDVS N-Caltech101
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
1 0.316 0.771 0.576 2.446 0.440 0.538
5 0.348 0.826 0.615 2.626 0.468 0.580
6 0.397 0.853 0.639 2.619 0.477 0.599
7 0.396 0.858 0.646 2.701 0.482 0.604
8 0.394 0.860 0.648 2.702 0.477 0.586
Few-shot Evaluation. We investigate the impact of training data scale on event-based understanding performance. As shown in Table VII, performance is closely tied to the dataset size, which aligns with the principle of Scaling Law [32]. When $n=5$, the model acquires more class-invariant knowledge, leading to more comprehensive perceptions. We also observe that with a small dataset ($n\leq 10$), our model achieves significant improvements in event analysis, demonstrating the superiority of our design. However, the enhancement in event-based understanding gradually diminishes when $n\geq 20$, as event-based knowledge learning approaches its limits within the constraints of the model’s parameters.
TABLE VII: Quantitative results about few-shot caption generation. “nn” denotes the sample number per category used in training.
Val set $\bm{n}$ Bleu-4 METEOR ROUGE-L CIDEr
N-ImageNet 1 0.058 0.103 0.308 0.398
5 0.081 0.122 0.345 0.576
10 0.092 0.131 0.354 0.673
15 0.094 0.133 0.365 0.699
20 0.099 0.138 0.363 0.742
HARDVS 1 0.216 0.210 0.541 1.244
5 0.343 0.278 0.607 2.017
10 0.356 0.299 0.625 2.177
15 0.378 0.314 0.628 2.457
20 0.405 0.318 0.641 2.721
Event Representation. As shown in Table VI, we observe a significant performance degradation when excluding the dynamic patch setting ($n_{\max}=1$). Lower values of $n_{\max}$, such as $n_{\max}=5$, result in coarse alignment between images and events due to insufficient semantic information. In contrast, increasing $n_{\max}$ to values like $n_{\max}=8$ does not further enhance EventVL’s understanding, as its category knowledge reaches inherent limits. Furthermore, larger values of $n_{\max}$ lead to higher CUDA memory consumption. To strike a balance between performance and resource efficiency, we select $n_{\max}=6$ for optimal temporal interaction and semantic extraction. We also investigate various levels of aggregation in the Adaptive Spatial Split process. As shown in Table VIII, results demonstrate that aggregating events at all levels yields the best performance, while omitting any level of representation leads to a effectiveness decline .
TABLE VIII: Different level of the event-data aggregation choice for event batch construction. “L1/L2/L3” denotes the Level-1/2/3.
L1 L2 L3 N-ImageNet HARDVS N-Caltech101
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
- - 0.332 0.784 0.593 1.923 0.443 0.559
- 0.356 0.805 0.610 2.457 0.454 0.565
- 0.360 0.802 0.613 2.459 0.450 0.568
0.397 0.853 0.639 2.619 0.477 0.599
TABLE IX: Various loss construction for training objective Equation 5. “Prior” denotes the image prior and $\mathcal{L}_{c}$ denotes the image-event embedding similarity loss.
Prior $\mathcal{L}_{c}$ N-ImageNet HARDVS N-Caltech101
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
- - 0.321 0.740 0.534 2.036 0.330 0.475
- 0.349 0.762 0.579 2.442 0.365 0.501
- 0.363 0.749 0.556 2.420 0.341 0.534
0.397 0.853 0.639 2.619 0.477 0.599
TABLE X: Different $\lambda_{1}$ and $\lambda_{2}$ setting for the training objective of Equation 5.
$\lambda_{1}$ $\lambda_{2}$ N-ImageNet HARDVS N-Caltech101
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
0.2 1 0.304 0.375 0.522 1.260 0.330 0.406
0.5 1 0.325 0.449 0.564 1.580 0.337 0.444
1 0.1 0.297 0.680 0.476 1.911 0.387 0.323
1 0.2 0.304 0.703 0.498 2.144 0.395 0.348
1 0.5 0.352 0.739 0.585 2.308 0.421 0.460
1 1 0.397 0.863 0.639 2.619 0.477 0.599
Optimization Setting. As shown in Table IX, we first analyze the main components of our training objectives. When using only cross-entropy for event-text alignment, model performance decreases by 9.7% in CIDEr on the N-ImageNet evaluation. Incorporating the image prior improves performance, likely by bridging the semantic gap between text and event data. The proposed $\mathcal{L}_{c}$ introduces an additional 5.8% improvement in CIDEr on N-ImageNet, further enhancing multimodal alignment. We also investigate the impact of varying weight settings in the optimization objectives of Equation 5. As shown in Table X, the results indicate that text-based alignment has a greater influence on the model’s understanding of event data. Lower values of $\lambda_{1}$ correspond to a more significant drop in performance, highlighting its importance in our optimization strategy.
VI Conclusion
In this work, we introduce EventVL, the first event-based Multimodal Large Language Model (MLLM) designed to achieve a comprehensive understanding of event stream data. By integrating an advanced event encoder that captures rich semantic information, a novel Event Spatiotemporal Representation for efficient feature aggregation, and a Dynamic Semantic Alignment mechanism for fine-grained feature alignment, EventVL significantly enhances the understanding and reasoning capabilities for event-based data. Additionally, we contribute a large-scale, high-quality annotated dataset with nearly 1.4 million event-image-text pairs, providing a robust foundation for future event-based research. We believe that EventVL sets a new direction for chat-centric event stream comprehension and lays the groundwork for future breakthroughs in event community.
Appendix A Related Work
Event-based Vision. The microsecond-level temporal resolution, high dynamic range (typically 120 dB, compared to the 60 dB of standard cameras), and power efficiency of event cameras represent a paradigm shift in imaging, surpassing traditional frame-based methods [1]. A wide range of event-based low-level imaging tasks have been explored, including recognition, perception, localization, and reconstruction. These span applications such as object recognition [33, 34], object detection [3, 35], optical flow estimation [36, 37], semantic segmentation [9], depth estimation [38], and object reconstruction [39, 40, 41, 42, 43], often leveraging techniques like NeRF and Gaussian Splatting. While event-based models have demonstrated success in these conventional perception tasks, the potential of event-based Multimodal Large Language Models (MLLMs) remains largely unexplored. In this work, we introduce a novel application: event captioning and description for comprehensive event stream understanding. This approach aims to address the challenges posed by the sparse, asynchronous, and high-temporal-resolution nature of event data, enabling the generation of explicit language descriptions. Such capabilities are particularly critical for safety-sensitive applications, including drone-based and in-vehicle perception systems.
Open-Source MLLMs. The emergence of open-source MLLMs [13, 44, 19, 18] has profoundly impacted the field of Artificial General Intelligence (AGI) by bolstering the ability to process both visual and textual data. In the past year, several open-source MLLMs have gained prominence such as MiniGPT-4 [45] and VisionLLM [46]. However, these models typically rely on traditional images or videos as input, which can result in the generation of low-quality images under extreme imaging conditions, thereby impairing comprehensive world understanding. Event cameras, with their distinct properties, can effectively address these challenges. Therefore, we are exploring the integration of event-based data into existing MLLMs to unlock the full semantic potential of event data.
Appendix B Data Engineering
A high-quality dataset is crucial for training a MLLM. Some datasets pairing images or videos with events [15, 16] offer only basic categorical text annotations, akin to ”This is a category”, which lack detail and variety. While certain studies suggest using coarse image-text data to address the limited diversity of image-event datasets like N-ImageNet, these annotations often have a weak connection to the images and include much extraneous information. Our experiments indicate that training with such datasets leads to a decline in performance as shown in Figure XI. Therefore, it is imperative to re-annotate this data to create a more precise and granular dataset, thereby enhancing the MLLM’s comprehension.
As shown in Figure 5, we mainly design a simple but effective generation pipeline for different domain datasets that contain static images, dynamic scenes and human motions. Motivated by previous works [29, 11], we chose the InternVL2-76B [17] as the foundation model for description and dialogue data generation, which has shown its superiority and surpasses several SOTAs such as GPT-4o or Claude 3.5 on various understanding benchmarks. Utilizing it to label our data can reduce multiple costs compared to commercial engines and bring higher-quality descriptions. Our captioning problem lists and detailed tasks are shown in Figures 6 and 8.
Coarse Generation. We first manually design the prompts and problem lists corresponding to various datasets, which can be referred to in our appendix. For image-event datasets (N-ImageNet, N-Caltech101), we first randomly sample questions from the problem list, then directly send images and event-based frames with questions into the InternVL2-76B for generating coarse answer. For video-event datasets (HARDVS, DSEC), it is inefficiency that utilizes all frames for generating description, which causes a large amount of computational consumption [9, 47]. To effectively extract long-term information from origin videos, we sampled uniformly $n$ frames. If the video contains less than $n$ frames, we adopt all frames. For balancing the cost and performance, we chose $n=14$. Then, the video, the instruct and event frames are sent to InternVL2-76B for getting detailed description.
Manual Check. The cost of manually checking every image/video-eevnt-text description is enormous. To address this, we develop sampling strategies based on the category classification of each dataset. For N-ImageNet, N-Caltech101, HARDVS, and DSEC, these datasets contain 1000 classes, 101 classes, 300 classes, and 20 scenes, respectively. We sample every coarse description per class or scene for evaluation. If the generated description is unsatisfactory, all images/videos from the corresponding class or scene are sent to the MLLM for caption regeneration using modified prompts and designated problems. If the regenerated captions still do not meet our quality requirements, we manually review and adjust the answers.
Finally, we construct the EventVL-Base dataset, comprising 1,393,444 meticulously annotated multimodal pairs with rich event descriptions. Additionally, we introduce EventVL-QA, a high-quality dataset containing 75,460 QA pairs designed to enhance model dialogue and comprehension capabilities. Specifically, EventVL-QA includes multiple QA pairs per category. For N-ImageNet, HARDVS, and DSEC, we define 10, 20, and 20 carefully curated sequential questions per category, respectively, to improve the model’s reasoning ability. The dataset scales are detailed in Table I and Table II, while additional dataset details can be found in the supplementary materials.
Appendix C Preliminaries
The original stream output of an event camera consists of a set of asynchronously generated events, where each event is represented by $(x,y,t,p)$. Here, $(x,y)$ denote the spatial coordinates, $t$ is the timestamp, and $p$ indicates the polarity of the intensity change, with +1+1 representing an increase in brightness and 1-1 a decrease. These events exhibit an exceptionally high dynamic range exceeding 120 dB and a temporal resolution equivalent to thousands of frames per second. However, directly processing event streams remains challenging [1]. To facilitate cross-modal alignment, previous works have commonly transformed event data into frame-like representations, such as grayscale images [12] or red-blue color maps [9, 7].
Appendix D Experiments
D-A Implementations
Fundamental Module. We selected the InternViT-300M-448px  [^2] https://huggingface.co/OpenGVLab/InternViT-300M-448px [17] as the foundational architecture for both the image and event encoders, due to its proven effectiveness across various image and video understanding benchmarks. This model consists of 24 attention blocks, as illustrated in Figure 7. The term “Drop” refers to the drop path, which is commonly applied in the main path of residual blocks. In our setup, $B$ represents the batch size, $C$ is the number of channels, and $H$ and $W$ denote the height and width of the image or event, respectively. $P$ is the patch number, and $C_{\phi}$ is the final output embedding channel size. We set $H=W=448$, $C_{\phi}=2048$, and $P=256$. We also leverage the pretrained weights of InternViT for initialization, enabling the transfer of spatial prior knowledge from images to events, which accelerates training and enhances data efficiency for EventVL. For the text decoder, we chose InternLM2-1.8B [48], which has demonstrated significant improvements in reasoning and coding tasks.
Figure 5: Illustration of our proposed data generation pipeline. By utilizing this simple pipeline, we propose two kinds of dataset for our model training – EventVL-Base and EventVL-QA datasets, which contains the rich description and high-quality multiple QA pairs.
Figure 6: Captioning Problem list for different kinds of datasets, which are utilized for EventVL-Base dataset. “Static Images” is for N-ImageNet [15] and N-Caltech101 [26], “Human Motions” is for [16] and “Drive Scenes” is for DSEC [25].
D-B Qualitative results
We also present additional examples to validate the superiority of our proposed EventVL framework, using only event data as input as shown in Figure 9 and Figure 10.
Figure 8: Illustration of our question lists prepared fro EventVL-QA datasets. Note that for DSEC [25] and HARDVS [16] scenes, we sample 20 problems per scene for the corresponding list randomly.
Figure 9: Qualitative results about processed zero-shot caption generation by utilizing N-ImageNet [15] val set from Event-Base dataset. Red color denotes model wrong description and orange color denotes the accurate object description. Underline text means redundant description.
Figure 10: Qualitative results about processed zero-shot caption generation by utilizing HARDVS [16] val set from Event-Base dataset. Red color denotes model wrong description and orange color denotes the accurate object description.
D-C Ablation Study
Dataset Quality Assessment. In this section, we first assess the effectiveness of our proposed dataset. As shown in Table XI, we test the untrained EventVL initialized from InternVL2-2B on all validation sets to evaluate the performance of current SOTA MLLM weights on the processed dataset. Compared to the results in Table XI, our method achieves 1.769 CIDEr (+46.3%) on the N-ImageNet validation set, significantly outperforming the untrained model (1.306 CIDEr). Experiments show that event representations in the red-blue color map confuse the MLLM, leading to a significant degradation in understanding. Furthermore, event data hampers the MLLM’s comprehension of images. These observations highlight the event-based understanding deficiencies in MLLMs and underscore the importance of our approach. Next, we investigate the enhancement of model training performance, as shown in Table XII. We conclude that training on a single-domain dataset positively impacts the corresponding validation sets, with only minor improvements on other validation sets. Our findings also suggest that training across multiple domains improves the model’s general event-based understanding, mitigating overfitting issues commonly seen in single-domain training. For instance, this approach results in nearly a 20% increase in CIDEr on the N-Caltech101 dataset. Moreover, incorporating scene data from DSEC enhances the model’s ability to capture broader contextual details, as evidenced by CIDEr gains on HARDVS when compared to training without DSEC. Finally, these improvements in event-based understanding significantly enhance the model’s overall multimodal comprehension.
TABLE XI: Exploration experiments of using the untraining EventVL initialized from InternVL2-2B [17]. “wo ev./im.” denotes that we don’t use the event or image modality as input.